OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through the point (4, -1) and is perpendicular to the line whose equation is 2x - y - 7 = 0?

OpenStudy (anonymous):

The slope intercept form of the given line is y=2x-7. The slope of this line is 2. To have a line perpendiciular the slope of the new line should be -1/2.

OpenStudy (anonymous):

y=mx+b (4,-1) -1 =-1/2(4)+b solve for b. Then rewrite using the slope and point b in slope intercept form.

OpenStudy (anonymous):

y=-1/2x+1

OpenStudy (anonymous):

okay thanks for explaining it ;)